# Model Predictive Control of the Neural Manifold

### v1

@@ 2.2 Dimensionality Reduction of Stimuli and Neural States
The control problem can now be expressed as given a latent embedding of the neural states and stimulus states ,
we seek to find an optimal set of latent inputs
with the dynamics model
@@
The SNN can then be stimulated with the decoded latent inputs
@@
2.3 Stimulus VAE
A VAE was trained to find a low dimensional representation of MNIST digit stimuli (sVAE). The sVAE encoder was composed of two single-channel convolutional layers and a single feedforward layer. Each of the three layers used a ReLU activation function. The size of the latent space was chosen to be 2. The sVAE decoder was symmetric to the architecture of the encoder with one key difference; a sigmoid activation function was used on the final output layer to ensure that the values of the reconstructed stimuli were between 0 and 1 (the bounds of all pixel values in the training images). See Supplementary Information T2 for layer hyperparameters.
The second half of the MNIST data set () was used for training and validation of the sVAE. Eighty percent of this data () was used for training and the remaining 20% for validation (). The sVAE was trained via gradient descent in Pytorch using the Adam optimizer.
A sequence of latent inputs for latent dynamics model identification was constructed using the validation set of images. Discrete points in the latent sVAE space were obtained by running k-means clustering (n=100) on the low dimensional embedding of the validation images. Half of the centers were used for the dynamics model training () and half for testing (). These discrete points were converted to a continuous time series by one of three methods: step function, fast interpolation, slow interpolation. The step function method took a sequence of centers and held each center constant for 500 time steps (ms). The fast and slow methods took the sequence and linearly interpolated values between each element in the sequence but at different time scales (200 ms for the fast and 1000 ms for the slow). Both the and sequences were 46.3 seconds long (46,300 time steps). The use of these three methods was to have a input sequence that could show how the latent states of the SNN responded to inputs changing at different time scales and frequencies. Figure 2 shows the latent input sequences and their sVAE decoded values.
2.4 Neural VAE
The SNN was stimulated using the sVAE-decoded and input sequences with the resulting spiking activity used to fit a VAE for the neural states (nVAE). A random sample of 20% of the neurons () in the SNN was used to build the model. These neurons were the only units in the SNN that were measured for the entirety of experiments I and II. Recall that in a typical extracellular recording, only a subset of the neurons are observable. The purpose of this random sampling was to mimic the incomplete information that would be obtained in a real recording.
where was chosen to be and .
The nVAE encoder and decoder were symmetric, with each having five feedforward layers. The smoothed state was z-scored when entering the first layer of the encoder. A small value was added to the denominator of the standardization since some time steps had near identical values. Each hidden layer used a ReLU activation function followed by a batch normalization layer. The size of the latent dimension was chosen to be 2. Because the addition of noise to the rLIF neurons resulted in temporal spiking jitter to the same input stimulus, the variance of the latent representation was constrained to be above 1. This was achieved by having the encoder learn the log of the variance instead of the variance directly and then applying the softplus function to the estimate. By using this minimum variance, it acted as a regularizing parameter to the noise present in the training data. See Supplementary information T3 for layer and training hyperparameters.
2.5 Latent Dynamics Model
A linear latent dynamics model of the form
was estimated using ridge regression with the L2-hyperparameter chosen via leave-one-out cross-validation. All model fitting was performed using the sklearn Python package. The dynamics model was fit with the data produced by the input sequence. Model performance was assessed by using an initial value of and forecasting for the entire training sequence length. At every time step, the known value of and the model’s previous prediction of was used to forecast the next value. An inadequate model would produce a time-series that was a poor approximation of the actual latent state trajectory produced by and could possibly even diverge due to compounding errors.
As an example, the forecasted values of the training data in Experiment I were a close fit to the actual training trajectories as measured with the product-moment correlation (, ). The possibility of overfitting was assessed by forecasting with the latent states elicited from . The resulting predicted trajectory was also similar to the actual latent trajectory (, ) indicating that the dynamics model would be useful for control with MPC. See Figure 4 for the forecasted latent trajectories of the training and testing data in Experiment I.
@@ 3 Results
3.1 Experiment I: MPC Can Force the Latent Dynamics on the Neural Manifold to Stay at set points
As a simple first experiment, the latent dynamics were controlled to follow a step function. This reference trajectory was composed of two set points in latent space, each held constant for 500 ms. The values were chosen by running k-means clustering on the latent state training data and using the resulting centroids as the set points. The controller had a predictive time horizon of 30 time steps and optimized the loss function
@@ 3.1 Experiment I: MPC Can Force the Latent Dynamics on the Neural Manifold to Stay at set points
No constraints other than the dynamics model were included. Due to the stochastic nature of the SNN responses, 50 independent trials of MPC were performed. The normalized mean square error (nMSE) was used to quantify the controller performance of the latent state, which normalizes mean square error by the difference between the maximum and minimum values. This was done for ease of comparing controller performance across dimensions with different scales. All MPC optimizations and implementations were performed using the do-mpc python package [35]. This package utilizes CasADi [36] and IPOPT [37] for interior-point optimization and automatic differentiation methods.
@@
The controller was able to achieve good performance, especially when considering the noise in the system and that only 20% of the neurons of the SNN were observable. Even though the noise that was added to all neuron inputs was Gaussian, the nonlinearities in the rLIF models propagate highly complex noise structures throughout the network. At the beginning of control and after the set point was switched, the system converged to the desired location in the latent space within 20–50 time steps and then remained there despite continual noise inputs that produced large perturbations (Figure 5A). When the set point was changed, there was a large excursion of the control stimulus within its latent space, and there were large fluctuations in the input throughout the experiment as the controller corrected for noise-induced perturbations.
@@
Because the latent space has only two dimensions, we can visualize the dynamics of the forecasting model as a vector field and the current and desired states of the system as points within this field. The dynamics can be decomposed into an autonomous component and the forcing from the stimulus , and as the combined system is linear, there is at most a single equilibrium. As seen in Figure 5B, at time points when the set point was constant, the controller chose stimuli such that the dynamics had a stable equilibrium close to the set point. During the transition to a new set point, the controller initially moved the equilibrium to a location beyond the set point but then settled to an input that created a fixed point at the desired location. This overshoot may have allowed the controller to more rapidly move the system to its desired location.
@@
As seen in Figure 5C, there was a qualitative change in the activity of the population when the reference point changed at 500 ms. Interestingly, while there was an initial drop in the average firing rate, it returned to a similar value as the system converged on the new set point. This indicates that the latent states are not just functions of the population firing rate, but are functions of particular firing patterns and activity from specific neurons. We can also examine the stimuli that were actually presented to the SNN during control. As seen in Figure 5D, the control stimuli all had digit-like characteristics, which undoubtedly reflects the fact that the sVAE was trained only on digits. Notice, however, that the image presented during the transition to the second reference point is a larger and more intense ‘2’ than the image presented after the system had converged.
@@
3.2 Experiment II: Distinct Reference Trajectories Produce Different Levels of Controller Performance
@@
The previous experiment revealed how the controller successfully used the stimulus-dependent dynamics of the forecasting model to drive the network to specific locations in the latent space. However, the forecasting model is only a linear approximation of the simulated network’s dynamics, which are nonlinear and of a much higher dimension. To examine how control could be used to probe the underlying dynamics of the network, it is not sufficient to characterize the start and end points of a trajectory, but instead what specific path was taken. For our next experiment, the reference trajectory was replaced with two time-varying functions (reference trajectory 1 and 2). Each of these reference trajectories had the same initial () and final values (), but took different paths through latent state space. Using a parameterized function of a circle passing through points and , trajectories 1 and 2 were the opposite arcs of the resulting circle. This ensured that both trajectories were of equal length through the latent state space. The observation model from Experiment I was used for this task (20% observability) and the values of and were the set points from the same experiment. Fifty control trials were performed for each of the reference trajectories with each trial having the same MPC hyperparameters from Experiment I.
As seen in Figure 6A, the controller successfully forced the network to take two different trajectories through the latent space, both of which were different from the trajectories the system used in Experiment I. However, the paths the system took often deviated substantially from the reference trajectory. For both trajectory 1 and 2, it appeared as though the system was attracted to specific regions of the state space and that there were other regions that the controller was unable to force the system to enter. These regions likely correspond to attractors and separatrices in the true dynamics of the SNN. Interestingly, individual trajectories appeared to show large oscillations in some regions of the state space, which could reflect rapid switching between stable equilibria or a limit cycle. It remains to be seen if a nonlinear forecasting model would enable the controller to force the system through repulsive regions of the state space better (see Discussion). There were differences between the two trajectories in the average state error (Figure 6B) and the variance of the latent inputs produced by the controller (Figure 6C). This indicated that more “power” in the latent space was needed to control the system to follow trajectory 2, even though the lengths of the two trajectories were the same.
@@
Examining the activity in the measurement space, we see obvious differences in the spiking behavior of the SNN (Figure 7A). The visual stimuli produced from decoding the latent inputs were initially similar between the two reference trajectories but then diverged as the trajectories separated (Figure 7B). This is consistent with the behavior of the SNN, which was trained to produce distinct responses to different digits; thus, different regions of the state space are likely to correspond to specific digits. One implication of this is that MPC of the latent dynamics could reveal if specific kinds of stimuli correspond to particular latent trajectories or if the activity in the latent space is driven by the differences in the stimulus at every time step (e.g. prediction error [38]).
3.3 Experiment III: The Proportion of Observable Neurons Affects Controller Performance
To investigate how robust the control strategy was to partial observation of the neural population, the procedure from Experiment I was performed on different subsets of neurons randomly sampled from the SNN. For each level of neuron observability, 10 independent ensembles were sampled from the SNN (). An nVAE was trained for each ensemble, but in order to ensure a fair comparison, the architecture was kept constant (except for the size of the input layer). As shown in Figure 8A, there was a tradeoff between the performance of the nVAE and the performance of the forecasting model. With a small number of neurons, (1% observability, 6 neurons), the nVAEs achieved almost perfect reconstruction for both training and testing data. However, the forecasting models performed very poorly, indicating that the stimulus-dependent dynamics of the full SNN could not be inferred from the behavior of just a few neurons. With larger proportions of the population observed, the performance of the nVAEs decreased, indicating that a latent space with only two dimensions was not sufficient to capture all the variance in the neural activity. There was a floor to this degradation of performance, and the reconstruction errors plateaued around 30% observability. This is consistent with a high level of correlation within the SNN. The performance of the forecasting models increased with the number of observed neurons. With 20% or more of the neurons included in the model, the dynamics of the SNN within the latent space could be predicted with a high level of accuracy.
The latent dynamics inferred at each level of observability were controlled to follow a step function of two set points obtained from k-means clustering, as in Experiment I. Ten trials of MPC were run for each of the models using the same previous loss function and hyperparameters. In practice, the hyperparameters of the controller would be tuned after fitting the forecasting model, but keeping them the same provides a fairer basis for comparison. Overall, controller performance was better with larger proportions of neurons observed (Figure 8B,C). Additionally, the variability of average nMSE across ensembles was typically higher when fewer neurons from the population were observed. This is unsurprising since there are many more possible ensembles to sample in the lower observation conditions which may differ in their controllability. However, it was still possible to sample bad ensembles when observing a large proportion of the neurons. The three highest levels of neuron observability all had an ensemble that produced higher average nMSE for a single latent state than the average of the nMSE distribution of the three lowest levels (Figure 8B).
@@ 4 Discussion
In this study, we demonstrated that MPC can be used to control the latent states of an artificial spiking neural network using low-dimensional, data-driven models of the dynamics. Despite the highly nonlinear dynamics of the SNN, the latent states could be fixed to certain set points or forced to follow arbitrary, time-varying reference trajectories. Control was possible even when only a limited proportion of the neurons was observed and there when there were unknown sources of noise in the network. Reducing the dimensionality of the neural activity and the stimuli used to control the network can therefore make MPC computationally tractable for high-dimensional neural systems. Based on this approach and this proof of principle, it should be possible to use MPC for real-time control of latent dynamics in sensory-driven biological networks using extracellular recordings. Achieving this level of control over neural networks would enable experimenters to probe the dynamics of neural circuits in ways that we have begun to outline here.
@@
In Experiment II, we found that the state of the network could be forced to take different trajectories by providing distinct sequences of latent inputs, but there were regions of the latent space where control was better or worse. One possibility is the latent dynamics model made better predictions for certain regions of latent state space, where the dynamics of the full network were more linear. The use of a nonlinear dynamics model may result in a better approximation of the latent vector field, but at the cost of complexity in both model fitting and MPC optimization. Another possibility is that the VAEs may have learned a dimensional transformation that preserved the distributional information in the high-dimensional space but that distorted the temporal relationships needed to infer the latent dynamics in certain regions [11]. By fitting the VAEs and latent dynamics models in two distinct phases, this may have produced dynamics models that were not optimal for the subspaces found by the VAEs. An alternative approach would be to simultaneously learn the dimensionality reduction and latent dynamics model in a single step, for example by extending a recurrent switching linear dynamical systems approach [13] to nonlinear dimensional reduction or by using structured variational autoencoders [39]. Although this would likely improve the performance of the controller, it would also introduce greater complexity when fitting the model, as additional hyperparameters would be needed to scale the influences of the VAE reconstruction, latent forecasting, and measurement space forecasting. Extensive work may be needed to identify the best approaches for balancing forecasting accuracy with computational efficiency for a given system and set of scientific questions.
@@
Recent years have seen the development of high-density silicon electrodes that can record extracellular spikes from hundreds to thousands of neurons simultaneously [40, 41], but this is still only a tiny fraction of the number of cells that participate in local neural circuits. In Experiment III, we found that the average performance of the forecasting model and the controller decreased as the proportion of neurons observed in the SNN decreased. However, these lower observation conditions also had higher variance in the distribution of average nMSE, indicating that good control over small ensembles is possible. It is also well-known that when observing a subset of states of a dynamical system, time-delay embedding the measurements gives information on the full dynamics [42]. In each of the experiments here, time-delay embedding was not used to fit the VAEs or latent dynamics models for simplicity. Although this did not appear to impact the controller performance for the higher observation percentage models, it may have resulted in poorer performance as the percentage decreased in value. It would be interesting to examine if using time delays would result in increased performance even when the number of observed neurons is very low. Other work has shown the success of using time-lagged autoencoders to find latent dynamics models [43], and methods exist to find optimal time delays and the number of embedding dimensions [44]. This would introduce additional complexity in practice however, because both the dimensionality reduction and dynamics model need to be fit quickly when recording from living neurons, otherwise the experiment may not be feasible in a laboratory setting due to cell death or electrode drift.
@@
The divide between the descriptive and generative perspectives on neural manifolds is related to the question of whether there is a low-dimensional dynamical system that emerges from the much larger and more complex dynamical system defined by the biophysics of intrinsic and synaptic currents in large populations of neurons. In this study, the true dynamics of the SNN were defined by a large, nonlinear dynamical system composed of hundreds of interconnected artificial neurons. The activity of this network could be represented in a low-dimensional neural manifold (Figure 3) with a clustered structure that is suggestive of attractor basins. The MPC framework developed in this study provided a method for experimentally probing the dynamics on this manifold to better understand its structure and function. Our results are consistent with the generative perspective in that a simple linear approximation of the dynamics in the latent space was sufficient to achieve a high level of control over the network. However, we also found that that there were regions of the state space where the latent dynamical model was not a good enough model of the underlying system to provide strong control, a result that can be interpreted as support of the descriptive perspective or as a source of insight for how to improve the latent model. If MPC can be applied in biological systems, it could provide a strong test of whether manifold activity is causal by enabling experimenters to see if specific organismal behaviors can be produced by controlling the latent neural dynamics. This would be an powerful tool in furthering the understanding of how complex behaviors and computations emerge from the structure of neural circuits and the dynamics of their activity.

### v2

@@ 2.3 Neural Manifold and Latent Dynamics Model Identification
We can now restate our original control problem as follows: given a latent encoding of the neural states
we seek to find a latent control input such that it forces to follow some desired trajectory . This control input can be found easily for PID by scaling the present and past errors by their corresponding gains in equation (5). For MPC, we find an optimal set of latent inputs
given the dynamics model
@@
and any additional constraints. For both control strategies, the AC can then be stimulated with the decoded latent inputs
@@
There are many ways to find the latent embedding of the neural activity and the resulting latent dynamics model. One approach is to identify the neural manifold by applying standard dimensionality reduction methods to recorded activity such as PCA and autoencoders [8] and then identify the LDM within this latent space using data-driven modeling with sequences and . However, learning the neural manifold separately from the dynamics can lead to suboptimal performance, because the latent space may not be well-aligned with the system’s underlying dynamics [12]. Several methods have been proposed to simultaneously learn latent dynamics and structure [35, 36, 37, 38]. A popular method is Latent Factor Analysis via Dynamical Systems (LFADS) and its related variants [39, 40], but these are based on recurrent neural networks (RNNs), which can be challenging to train and interpret. Instead, for simplicity we use a VAE-based reduction of the neural states and implement a dynamics model structure similar to LFADS but without hidden recurrent components. Our purpose here is not to innovate LDM architecture design, but to demonstrate how dimensionality reduction and data-driven dynamical systems modeling can be used to control the neural manifold. We suspect that the best model for neural manifold control will depend on the properties of the system, control apparatus used, and desired behavior.
Our approach consists of two stages: first, we pretrain a VAE to learn a structured latent representation of neural activity (nVAE). Then, we incorporate the pretrained nVAE and the sVAE into a larger dynamics model that learns linear dynamics within the latent space. While the parameters of the nVAE are free to vary during LDM training, the parameters of the sVAE are frozen, under the simplifying assumption that the latent structure of the stimulus is independent of the neural manifold’s structure, though the reverse is not necessarily true.
The nVAE was pretrained using the smoothed spikes to start the nVAE encoder and decoder at a good point for modeling the latent dynamics. The dataset had training, validation, and testing partitions corresponding to the stimulus set that produced them. During pretraining, the nVAE loss was given by
which had a corresponding structure to the loss function for the sVAE. The latent dimension of was chosen to be two. During pretraining, the learning rate was progressively reduced using a Lambda learning rate scheduler with a decay factor of 0.99. See Appendix 4 and 5 for nVAE architecture and pretraining hyperparamters.
After pretraining the nVAE for 20 epochs with , the full LDM was given by equations
The fine-tuning of the nVAE encoder and decoder as well as estimating unknown matrices and were found by minimizing the loss function
where the forecasting loss term ensures that the learned latent representations evolve consistently with the observed neural dynamics. Latent forecasts were obtained by encoding the observed spike behavior at time step as the initial value and using the latent input sequence to auto-regressively predict timesteps into the future. The model was trained until the loss from validation data and did not improve after 5 epochs. The learning rate was also reduced with a Lambda learning rate scheduler with a decay factor of 0.99. See Appendix 6 for LDM fine-tuning hyperparameters.
LDM performance was assessed by using an initial value of and forecasting for the entire training sequence length (n = 36,799). At every time step, the known value of and the model’s previous prediction of was used to forecast the next value (i.e., open-loop forecasting). There was a high correlation between the open-loop forecast and actual latent trajectory (, ) indicating that the dynamics model would be useful for control with MPC. See Figure 3 for LDM forecasting performance. While PID does not require an explicit model, we compare both PID and MPC within the same learned latent space parameterized by the LDM to evaluate their relative performance.
@@ 3 Results
3.1 Experiment I: Both PID and MPC Can Force the Latent Dynamics on the Neural Manifold to Stay at Set Points
As a simple first experiment, the latent dynamics on the AC neural manifold were controlled to follow a step function. This reference trajectory was composed of two set points in latent space, each held constant for 500 ms. The values were chosen by running k-means clustering on the latent state testing data and using the resulting centroids as the set points.
The MPC controller had a predictive time horizon of 10 time steps and optimized the loss function
@@
The values in the loss function matrices were hand-tuned and required fewer than a dozen runs to produce good performance. No constraints other than the dynamics model were included.
Because the latent control inputs were two-dimensional, the PID gains in equation (5) were all 22 matrices. Due to the difficulty in hand-tuning all the PID gains, a grid search of the PID gains was performed across each dimension with and set to . There were 6 candidate values for each element of : . Using the values that resulted in the lowest error in state tracking, we then manually tuned the value of and . The values of the PID gain matrices used in all experiments were
@@
Due to the stochastic nature of the AC responses, 50 independent trials of MPC and PID were performed. All MPC optimizations and implementations were performed using the do-mpc python package [41]. This package utilizes CasADi [42] and IPOPT [43] for interior-point optimization and automatic differentiation methods.
@@
Both controllers were able to achieve good performance, especially when considering the noise in the system and that only 20% of the neurons of the AC were observable. Even though the noise that was added to all neuron inputs was Gaussian, the nonlinearities in the rLIF models propagate highly complex noise structures throughout the network. At the beginning of control and after the set point was switched, the system converged to the desired location in the latent space within 20–50 time steps and then largely remained there despite continual noise inputs (Figure 4A). However, the PID control input had more variability between trials, suggesting that MPC was better able to account for the unknown full-state system dynamics and sources of noise.
@@
Because the latent space only had two dimensions, we can visualize the dynamics of the forecasting model as a vector field and the current and desired states of the system as points within this field. The dynamics can be decomposed into an autonomous component and the forcing from the stimulus , and as the combined system is linear, there is at most a single equilibrium. As seen in Figure 4C, at time points when the set point was constant, both controllers chose stimuli such that the dynamics had a stable equilibrium close to the set point. While the control input was highly consistent between trials for MPC, there was greater variability for the PID control inputs. In an experimental setting, this would result in less ability to distinguish the effects of the latent input from neural manifold dynamics.
@@
As seen in Figure 4B, there was a qualitative change in the activity of the population when the reference point changed at 500 ms for both controllers. We can also examine the stimuli that were actually presented to the AC during control. As seen in Figure 4C, nearly all the control stimuli had digit-like characteristics, which undoubtedly reflects the fact that the sVAE was trained only on digits.
3.2 Experiment II: MPC Hyperparameters Generalize Better to New Experimental Conditions Compared to PID
To investigate how generalizable the controller hyperparameters were to different experimental conditions, the procedure from Experiment I was performed on different subsets of neurons randomly sampled from the same AC. If the controllers from Experiment I did not generalize to other samples, there would be considerable time cost to performing this experiment in a laboratory setting. For each level of neuron observability, 10 independent ensembles were sampled from the SNN (n = 80). A LDM was trained for each ensemble, but in order to ensure a fair comparison, the architecture was kept constant (except for the size of the input layer). The nVAE and LDM training hyperparameters were kept the same as in Experiment I. The normalized mean square error (nMSE) was used to quantify the controller performance of the latent state, which normalizes mean square error by the difference between the maximum and minimum values. This was done for ease of comparing controller performance across dimensions with different scales.
As shown in Figure 5A, there was a positive, monotonic relationship between the proportion of neurons observed and performance of the LDM. Both the forecasting and nVAE decoding performances were low at only 1% observability, and there was a steep increase in performance that saturated at 20% observability. It is likely this saturation is due to the architecture design choices being tuned at the 20% level in Experiment I, but also indicates that there are diminishing returns to including more neurons in a recording session. To assess controller performance, the latent dynamics inferred at each level of observability were controlled to follow a step function of two set points obtained from k-means clustering, as in Experiment I. Fifty trials were run for each of the ensembles.
Shown in Figure 5B, the MPC hyperparameters from Experiment I did not generalize well to the 1% observation level, likely because of the poor LDM performance. However, as the observation percentage went up, the nMSE in MPC tracking error decreased, indicating that the hyperparameters generalized to these conditions. In contrast, the PID controller did not generalize well and had higher variability in average nMSE. As seen in 5B, except in the 1% observation condition, MPC had superior performance. Not only was the latent state tracking of PID even nosier than MPC, but the PID controller could catastrophically fail, producing a controlled trajectory far from the desired set point. As seen in 5C, the distribution of the average root mean square (RMS) of the control input had greater variability compared to MPC. This variability in control input suggests that MPC provides a more stable mapping between inputs and latent states, making it better suited for understanding the neural manifold compared to PID. If the goal is to design experiments that associate specific inputs with specific latent states, the oscillatory and inconsistent control signals of PID (as seen in 5E) would make this challenging. In contrast, MPC’s more stable and consistent control inputs lead to outputs that are easier to interpret in latent space. Given these results, we conclude that control over neural manifold dynamics is better achieved with MPC as opposed to PID due to its robustness to unknown noise and generalizable hyperparameters.
@@ 3.2 Experiment II: MPC Hyperparameters Generalize Better to New Experimental Conditions Compared to PID
3.3 Experiment III: Distinct Reference Trajectories Produce Different Levels of Controller Performance
@@
The previous experiments revealed how MPC used the stimulus-dependent dynamics of the forecasting model to drive the network to specific locations in the latent space. However, the forecasting model is only a linear approximation of the simulated network’s dynamics, which are nonlinear and of a much higher dimension. To examine how MPC could be used to probe the underlying dynamics of the network, it is not sufficient to characterize the start and end points of a trajectory, but instead what specific path was taken. In this experiment, the reference trajectory was replaced with two time-varying functions (reference trajectory 1 and 2). Each of these reference trajectories had the same initial () and final values (), but took different paths through latent state space. Using a parameterized function of a circle passing through points and , trajectories 1 and 2 were the opposite arcs of the resulting circle. This ensured that both trajectories were of equal length through the latent state space. The observation model from Experiment I was used for this task (20% observability) and the values of and were the set points from the same experiment. Fifty control trials were performed for each of the reference trajectories, with each trial having the same MPC hyperparameters from Experiments I and II.
@@
As seen in Figure 6A, the controller successfully forced the network to take two different trajectories through the latent space, both of which were different from the trajectories the system used in Experiment I. There were clear differences between the two trajectories in the RMS of the latent inputs produced by the controller (Figure 6C). In both cases, there was more “power” required for than in , but the difference in RMS was much larger for trajectory 2. Examining the activity in the measurement space, there were obvious differences in the spiking behavior of the AC (Figure 7A). The visual stimuli produced from decoding the latent inputs were initially similar between the two reference trajectories but then diverged as the trajectories separated (Figure 7B). This is consistent with the behavior of the AC, which was trained to produce distinct responses to different digits; thus, different regions of the state space are likely to correspond to specific digits. One implication of this is that MPC of the latent dynamics could reveal if specific kinds of stimuli correspond to particular latent trajectories or if the activity in the latent space is driven by the differences in the stimulus at every time step (e.g. prediction error [44]).
@@ 4 Discussion
In this study, we demonstrated that both PID and MPC can be used to control the latent states of an artificial spiking neural network. MPC had superior performance, which we attribute to its anticipatory nature and explicit use of data-driven models of the latent dynamics. Despite the highly nonlinear dynamics of the AC, the latent states could be fixed to certain set points or forced to follow arbitrary, time-varying reference trajectories. Control was possible even when only a limited proportion of the neurons was observed and there when there were unknown sources of noise in the network. Reducing the dimensionality of the neural activity and the stimuli used to control the network can therefore make MPC computationally tractable for high-dimensional neural systems. Although this may seem like a convenient trick that only works in our artificial example, we hypothesize this method of finding control inputs in the latent space can be extended to other experimental conditions. For example, naturalistic stimuli exhibit both linear and nonlinear correlations that can be used for dimensionality reduction in the stimulus space [45]. Based on this approach and this proof of principle, it should be possible to use MPC for real-time control of latent dynamics in sensory-driven biological networks using extracellular recordings. Achieving this level of control over neural networks would enable experimenters to probe the dynamics of neural circuits in novel ways. For example, perceptual decision boundaries have been modeled as attractor states in the neural manifold [46]. By controlling this latent activity along specific trajectories, researchers could more precisely investigate the visual features of the control stimulus that drive decision-making. Additionally, the framework we provide here can be extended to other modalities of stimulation. For example, instead of using external visual stimuli as , complex spatial patterns using two-photon holographic optogenetic stimulation could be used to control latent neural activity [47]. This could allow for exciting new experimental designs where the relationship between organism behavior and latent neural activity could be more deeply explored.
In Experiment II, we assessed the generalizability of MPC and PID controller hyperparameters by applying them to different ensembles of neurons randomly sampled from the same AC. Across all but the 1% observation condition, MPC demonstrated significantly better state tracking performance than PID, despite both controllers being tuned to a single ensemble in the 20% observation condition. Additionally, PID exhibited greater variability in its latent control inputs, which would make it more challenging to interpret the relationship between neural manifold activity and stimulus structure. One potential concern with our simulation is that recording even 1% of the neurons in a real neural circuit would be an exceptionally high percentage. However, we hypothesize that the poor performance of MPC in the 1% condition was primarily due to deficiencies in the learned dynamics model rather than the proportion of neurons recorded. As shown in Figure 5B, there were some ensembles that still resulted in very good control performance. While a detailed analysis of this issue is beyond the scope of the current study, future research should investigate how well MPC and other control systems perform in simulated circuits with different connectivity patterns and redundancy in neural responses.
@@
In Experiment III, we found that the state of the network could be forced to take different trajectories by providing distinct sequences of latent inputs, but there were regions of the latent space where control was better or worse. For both trajectories, it appeared as though the system was attracted to specific regions of the state space, and that there were other regions the controller was unable to force the system to enter. Interestingly, individual trajectories appeared to show large oscillations in some regions of the state space, which could reflect rapid switching between stable equilibria or a limit cycle. One explanation is that these regions correspond to attractors and separatrices in the true dynamics of the AC. The potential sources of these oscillations may have been obfuscated by our use of only two latent dimensions in all of our experiments. While this was done for ease of visualization and interpretation, a more rigorous analysis to determine the dimensionality of the latent space [48, 49] may have resulted in better forecasting and control. For example, the oscillatory behavior seen in some regions of the latent state space in Figure 6 may have been from a saddle point that would have been identified if we had used three or more latent dimensions. Another possibility is that the LDM made better predictions for certain regions of latent state space, where the dynamics of the full network were more linear. The use of a nonlinear dynamics model such as hybrid/switching linear dynamics [14, 50], time-varying dynamics [51], or structured variational autoencoders [52] may result in a better approximation of the latent vector field. Although this would likely improve the performance of the controller, it would also introduce greater complexity when fitting the model and in the MPC optimization. However, recent work for finding linear representations of nonlinear dynamical systems using Koopman theory offers promising evidence that a linear modeling approach can be effective for controlling neural manifold dynamics [12, 53]. Extensive work may be needed to identify the best approaches for balancing forecasting accuracy with computational efficiency for a given system and set of scientific questions.
An important practical consideration is how to decide if either controller is good enough for a particular application. We demonstrated that MPC outperforms PID control for an artificial network with strong sensory inputs, but the level of control MPC provides might still be lower than what would be needed to fully investigate the structure and function of the neural manifold in a biological circuit. The answer will surely depend on the specific system being studied and the biological questions, but we can suggest two heuristic criteria for evaluating whether a controller is sufficiently effective. First, the controller should produce smaller errors and variability with respect to a target neural trajectory than what is observed across repeated presentations of the original stimulus. In sensory experiments, there is always some variability in the neural or behavioral responses across trials where the same stimulus is presented. Controller-elicited responses should have a variability below the level of the natural sensory responses and regularly result in the desired trajectory. Achieving this would not only demonstrate control efficacy but also offer new insights into the sources of neural and behavioral variability. Second, an effective controller should be capable of identifying a stimulus or stimulus sequence that drives the system toward a desired behavioral or neural state that was not present in the training data used to fit the dynamics model. Success in this regard would indicate meaningful generalization of the latent dynamics and suggest a deeper understanding of the system’s structure. It is also important to note that the superior performance of MPC comes with increased computational complexity. MPC requires solving an optimization problem at each time step, which can lead to delayed control signals in real-time applications. This added overhead may limit its applicability when ultra-fast control is required or when computational resources are constrained. In contrast, PID control is significantly simpler and faster to implement, making it a more practical choice in scenarios where some loss in control accuracy is acceptable in exchange for speed and efficiency.
@@
Recent years have seen the development of high-density silicon electrodes that can record extracellular spikes from hundreds to thousands of neurons simultaneously [54, 55], but this is still only a tiny fraction of the number of cells that participate in local neural circuits. In Experiment II, we found that the average performance of the forecasting model and the controller decreased as the proportion of neurons observed in the SNN decreased. However, these lower observation conditions also had higher variance in the distribution of average nMSE, indicating that good control over small ensembles is possible. It is also well-known that when observing a subset of states of a dynamical system, time-delay embedding the measurements gives information on the full dynamics [56]. In each of the experiments here, time-delay embedding was not used to fit the VAEs or LDMs for simplicity. Although this did not appear to impact the controller performance for the higher observation percentage models, it may have been detrimental to performance as the percentage decreased. It would be interesting to examine if using time delays would result in increased performance even when the number of observed neurons is very low. Other work has successfully used time-lagged autoencoders to find LDMs [57], and methods exist to find optimal time delays and the number of embedding dimensions [58]. However, these methods may be too complex and computationally demanding to be used in practice, as recording time is often limited in vivo by electrode drift.
@@
The divide between the descriptive and generative perspectives on neural manifolds is related to the question of whether there is a low-dimensional dynamical system that emerges from the much larger and more complex dynamical system defined by the biophysics of intrinsic and synaptic currents in large populations of neurons. In this study, the true dynamics of the AC were defined by a large, nonlinear dynamical system composed of hundreds of interconnected artificial neurons. We were able to represent the activity of this network in a low-dimensional neural manifold (Figure 3) with a clustered structure that is suggestive of attractor basins. The MPC framework developed in this study provided a method for experimentally probing the dynamics on this manifold to better understand its structure and function. Our results are consistent with the generative perspective in that a simple linear approximation of the dynamics in the latent space was sufficient to achieve a high level of control over the network. However, we also found that that there were regions of the state space where the LDM was not a good enough model of the underlying system to provide strong control, a result that can be interpreted as support of the descriptive perspective or as a source of insight for how to improve the latent model. If MPC can be applied in biological systems, it could provide a strong test of whether manifold activity is causal by enabling experimenters to see if specific organismal behaviors can be produced by controlling the latent neural dynamics. This would be an powerful tool in furthering the understanding of how complex behaviors and computations emerge from the structure of neural circuits and the dynamics of their activity. Neural dynamics models as suggested in [35] would be particularly suited for these kinds of tasks since they are able to identify the dynamics directly related to behavior.